# Axis Tour: Word Tour Determines the Order of Axes in ICA-transformed Embeddings

### v1

@@ Acknowledgements
We would like to thank Momose Oyama for the discussion. This work was supported by JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2123. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3.
Appendix A Dimensionality reduction
Figure 5: Projection of the subspace spanned by three consecutive axes in Fig. 1 into a one-dimensional space. Each word is assigned the color of the axis with the highest value. The projection direction is in the direction representing the subspace. For visualization, we randomly sampled 10,000 words, excluding the top 5 words on each axis.
@@ A.1 Definition of $\mathbf{f}_{r}$
Skewness as weight.
In (3), a vector $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for dimensionality reduction is defined by the skewness $\gamma_{\ell}\in\mathbb{R}_{\geq 0}$ of the $\ell$-th axis of $\mathbf{T}$. In particular, $f_{r}^{(\ell)}$, which corresponds to the weight of the $\ell$-th axis in the projection, is proportional to $\gamma_{\ell}^{\alpha}(\alpha\in\mathbb{R}_{\geq 0})$. This is based on the assumption that the axis becomes more meaningful as the skewness increases [^12] While not specific to ICA-transformed embeddings, it is known from a study of sparse coding for language models that skewness correlates with interpretability Cunningham et al. (2023).. Higher-order statistics such as skewness are known to be sensitive to outliers Hyvärinen et al. (2001b), so we mitigate this effect by raising $\gamma_{\ell}$ to the power of $\alpha$. Since skewness is a third-order moment, we treat $\alpha=1/3$ as the default value. $\alpha=0$ is a uniform weight, while $\alpha=1$ is the original $\gamma_{\ell}$. Next, we will explain why $\mathbf{f}_{r}$ is normalized.
Reason for Normalizing $\mathbf{f}_{r}$.
@@
A.2 Projection from subspace to one-dimensional space
@@
Appendix B Details of experiment in Section 5
@@
Table 2: Dimensionality reduction performance for $p$-dimensional embeddings. Rand. stands for Random Order, Skew. for Skewness Sort, and Tour. for Axis Tour. The values in the table correspond to top 1 accuracy for analogy tasks, Spearman’s rank correlation for word similarity tasks, and purity for categorization tasks. Note that at $p=300$, all embeddings give the same results.
Figure 6: Dimensionality reduction performance for the PCA-transformed embeddings and the Axis Tour embeddings whose $\alpha=0,1/3,1/2,1$. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks.
The embedding of the word $w_{i}$ is denoted by $\mathbf{w}_{i}\in\mathbb{R}^{d}$.
@@
We used the Google Analogy Test Set Mikolov et al. (2013a), which contains 14 types of analogy tasks, and the Microsoft Research Syntactic Analogies Dataset Mikolov et al. (2013b), which contains 16 types of analogy tasks. In the analogy tasks, the quality of the embeddings is evaluated by inferring $w_{4}$ to which $w_{3}$ corresponds if $w_{1}$ corresponds to $w_{2}$. We compute the vector $\mathbf{w}_{2}-\mathbf{w}_{1}+\mathbf{w}_{3}$ and see if the closest embedding is $\mathbf{w}_{4}$ (top1 accuracy).
@@ Results.
Note that the experimental results are the same for all embeddings for $p=300$. First, as we saw in Appendix A.1, when $p=d\,(=300$), the matrix $\mathbf{TF}$ (i.e., the projected $p$-dimensional embeddings) is equal to the matrix $\mathbf{T}$ (i.e., the $d$-dimensional Axis Tour embeddings). Then, by definition, Axis Tour, Random Order, and Skewness Sort are the embeddings obtained by reordering the axes of the ICA-transformed embeddings and flipping their signs as needed. Thus, these three can be seen as the embeddings obtained by applying an orthogonal matrix to the ICA-transformed embeddings. Since the ICA-transformed embeddings are derived from the PCA-transformed embeddings by applying an orthogonal matrix [^15] Refer to the previous work for the relationship between PCA and ICA Yamagiwa et al. (2023)., cosine similarity and Euclidean distance remain unchanged for PCA, Random Order, Skewness Sort, and Axis Tour, leading to identical results in downstream tasks.
Appendix C Additional experiments
C.1 Comparison of $\alpha$

### v2

@@ Acknowledgements
We would like to thank Momose Oyama for the discussion. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110.
Appendix A Details of scatterplots from two-dimensional projection
This section explains the two-dimensional projection method used for the scatterplots in Fig. 1. We then present similar scatterplots for the three examples in Table 1. Finally, we define the metrics used in Section 1 to evaluate the quality of the scatterplots.
A.1 Scatterplot drawing method
We will explain the scatterplot drawing method using the Axis Tour embeddings.
First, we define a set of axis indices for projection. Let $I:=\{a,\cdots,b\}\,(a,b\in[d],a<b)$ be a consecutive interval of indices. The number of indices in $I$ is $|I|=b-a+1$. For example, in Fig. 1, $a=86$ and $b=94$.
Next, we define a matrix by extracting the axes of $\hat{\mathbf{T}}$ corresponding to $I$, where $\hat{\mathbf{T}}$ is the matrix obtained by normalizing the matrix $\mathbf{T}$. We denote this extracted matrix as $\hat{\mathbf{T}}_{I}\in\mathbb{R}^{n\times|I|}$.
[…]
$\mathbf{P}_{I}:=\begin{bmatrix}\varphi_{a}^{\top}\\ \vdots\\ \varphi_{b}^{\top}\end{bmatrix}\in\mathbb{R}^{|I|\times 2},$ (6)
where
$\varphi_{\ell}:=(\cos{\theta_{\ell}},\sin{\theta_{\ell}})^{\top}\in\mathbb{R}^{2},$ (7)
where
$\theta_{\ell}:=\frac{(\ell-a)\pi}{b-a}.$ (8)
Then we get the two-dimensional projection as $\mathbf{Q}_{I}:=\hat{\mathbf{T}}_{I}\mathbf{P}_{I}\in\mathbb{R}^{n\times 2}$. In $\mathbf{Q}_{I}$, the $\ell$th axis of $\hat{\mathbf{T}}$ is projected along the direction of $\varphi_{\ell}$.
We denote the $i$th word embedding of $\mathbf{Q}_{I}$ by $\mathbf{q}_{i}=(q_{i}^{x},q_{i}^{y})^{\top}\in\mathbb{R}^{2}$. When we plot the scatterplot of $\mathbf{Q}_{I}$, we do not plot the $i$th word embedding if $q_{i}^{y}<0$ for visual clarity, and show the top five words for each axis. The indices of these top words equal to the following index set $\text{Show}_{I}$ defined with $\text{Top}_{k}^{\ell}$ from section 4.1:
$\text{Show}_{I}:=\{i\in[n]\mid q_{i}^{y}\geq 0\}\cap\bigcup_{\ell\in I}\text{Top}_{5}^{\ell}.$ (9)
Similarly, we can apply the same procedure to the Skewness Sort embeddings and obtain the two-dimensional scatterplot.
A.2 Scatterplots of Table 1
(a) The 23rd axis to the 31st axis
(b) The 101st axis to the 109th axis
(c) The 237th axis to the 245th axis
Figure 5: The scatterplots of the two-dimensional projections for the axes of the Axis Tour embeddings in Table 1. Similar to Fig. 1, we used the procedure described in Appendix A.1.
Fig. 5 shows the scatterplots of the two-dimensional projections for the axes of the Axis Tour embeddings in Table 1, using the procedure described in Appendix A.1. Similar to Fig. 1, it is evident that the top words of the axes of the Axis Tour embeddings are farther from the origin than those of the Skewness Sort, and the meanings of the adjacent axes change continuously.
A.3 Evaluation metrics for scatterplots
In Section 1, we compared the quality of the scatterplots for Axis Tour and Skewness Sort by calculating the average distance of the top words from the origin. In this section, we first explain this metric and then, based on (4), define a new metric derived from the average of the cosine similarities between adjacent axis embeddings. We then compare these metrics for the scatterplots in Figs. 1 and 5. Similar to Appendix A.1, we use the Axis Tour embeddings to explain these metrics.
For the two-dimensional projection $\mathbf{Q}_{I}$ of the Axis Tour embeddings, we define the average distance $d_{I}$ of the top words from the origin, using the index set $\text{Show}_{I}$ in (9), as follows:
$d_{I}:=\frac{1}{|\text{Show}_{I}|}\sum_{i\in\text{Show}_{I}}\|\mathbf{q}_{i}\|$ (10)
A larger $d_{I}$ indicates that the scatterplot more accurately reflects the spatial distribution of the original embeddings, since the top words are positioned far from the origin.
We also define the average cosine similarity between adjacent axis embeddings, $c_{I}$, for the interval $I$ as follows:
$c_{I}:=\frac{1}{b-a}\sum_{\ell=a}^{b-1}\cos{(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})}$ (11)
As we saw in (4), since Axis Tour optimizes the order of the axes to maximize the sum of the cosine similarities, the value of $c_{I}$ reflects the semantic continuity of the axes in the scatterplot. It is important to note that $I$ represents a subinterval of $d$ indices, so we do not include the term of $\cos{(\mathbf{v}_{a},\mathbf{v}_{b})}$ in (11).
Both $d_{I}$ and $c_{I}$ can also be calculated in a similar manner for Skewness Sort. Table 2 shows the values of $d_{I}$ and $c_{I}$ for Figs. 1 and 5. In these examples, Axis Tour shows higher values for both $d_{I}$ and $c_{I}$ compared to Skewness Sort, indicating better projection quality.
In addition to the results for several subintervals in Table 2, we also see the semantic continuity of all axes. If $I=[d]$ in (11), we get $c_{I}=c_{[d]}$ as follows
$c_{[d]}=\frac{1}{d}\left\{\cos{(\mathbf{v}_{1},\mathbf{v}_{d})}+\sum_{\ell=1}^{d-1}\cos{(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})}\right\}$ (12)
where the term of $\cos{(\mathbf{v}_{1},\mathbf{v}_{d})}$ is added according to (4). The value of $c_{[d]}$ is equal to the average of the histogram in Fig. 2, so the value for Axis Tour is $0.244$ and for Skewness Sort is $0.017$, thus confirming the semantic continuity across all axes.
Axis Tour Skewness Sort
Fig. $d_{I}$ $c_{I}$ $d_{I}$ $c_{I}$
1 0.76 0.27 0.61 0.18
5 0.69 0.28 0.61 0.10
5 0.80 0.30 0.62 0.15
5 0.86 0.38 0.76 0.23
Table 2: The values of $d_{I}$ and $c_{I}$ for Figs. 1 and 5.
Appendix B Dimensionality reduction
@@ B.1 Definition of $\mathbf{f}_{r}$
B.1.1 Skewness as weight
In (3), a vector $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for dimensionality reduction is defined by the skewness $\gamma_{\ell}\in\mathbb{R}_{\geq 0}$ of the $\ell$th axis of $\mathbf{T}$. In particular, $f_{r}^{(\ell)}$, which corresponds to the weight of the $\ell$th axis in the projection, is proportional to $\gamma_{\ell}^{\alpha}\,(\alpha\in\mathbb{R}_{\geq 0})$.
This is based on the assumption that the axis becomes more meaningful as the skewness increases [^14] While not specific to ICA-transformed embeddings, it is known from a study of sparse coding for language models that skewness correlates with interpretability Cunningham et al. (2023).. Higher-order statistics such as skewness are known to be sensitive to outliers Hyvärinen et al. (2001b), so we mitigate this effect by raising $\gamma_{\ell}$ to the power of $\alpha$. Since skewness is a third-order moment, we treat $\alpha=1/3$ as the default value. $\alpha=0$ is a uniform weight, while $\alpha=1$ is the original $\gamma_{\ell}$. Next, we will explain why $\mathbf{f}_{r}$ is normalized.
B.1.2 Reason for Normalizing $\mathbf{f}_{r}$
@@
B.2 Projection from subspace to one-dimensional space
Figure 6: Projection of the subspace spanned by three consecutive axes in Fig. 1 into a one-dimensional space. Each word is assigned the color of the axis with the highest value. The projection direction is in the direction representing the subspace. For visualization, we randomly sampled 10,000 words, excluding the top five words on each axis.
@@ B.2 Projection from subspace to one-dimensional space
Appendix C Distribution of cosine similarity
Let us consider two random vectors $X=(X_{1},\ldots,X_{d}),Y=(Y_{1},\ldots,Y_{d})\in\mathbb{R}^{d}$ with elements of mean zero $\mathbb{E}(X_{\ell})=\mathbb{E}(Y_{\ell})=0$ and finite variance $\mathbb{E}(X_{\ell}^{2})=\sigma_{X}^{2}$, $\mathbb{E}(Y_{\ell}^{2})=\sigma_{Y}^{2}$. We assume that the elements $X_{1},\ldots,X_{d}$ and $Y_{1},\ldots,Y_{d}$ are independent, and the sequence $X_{1}Y_{1},\ldots,X_{d}Y_{d}$ satisfies Lindeberg’s condition Billingsley (1995); for $Z_{\ell}=X_{\ell}Y_{\ell}/\sigma_{X}\sigma_{Y}$, $\forall\epsilon>0,\lim_{d\to\infty}d^{-1}\sum_{\ell=1}^{d}\mathbb{E}\Bigl{[}Z_{\ell}^{2}1(|Z_{\ell}|>\epsilon\sqrt{d})\Bigr{]}=0$. Lindeberg’s condition means no one $X_{\ell}Y_{\ell}$ dominates the inner product $\langle X,Y\rangle=\sum_{\ell=1}^{d}X_{\ell}Y_{\ell}$, and it is satisfied, for example, when all the elements follow the identical distribution.
Then, for sufficiently large $d$, the cosine similarity $\cos(X,Y)$ asymptotically follows $\mathcal{N}(0,1/d)$, the normal distribution with mean 0 and variance $1/d$. In other words, for sufficiently large $d$,
$\sqrt{d}\cos(X,Y)\sim\mathcal{N}(0,1).$ (14)
This is easily shown as follows. First note that $\mathbb{E}(X_{\ell}Y_{\ell})=\mathbb{E}(X_{\ell})\mathbb{E}(Y_{\ell})=0$, $\mathbb{E}(X_{\ell}^{2}Y_{\ell}^{2})=\mathbb{E}(X_{\ell}^{2})\mathbb{E}(Y_{\ell}^{2})=\sigma_{X}^{2}\sigma_{Y}^{2}$. Thus the inner product, if scaled by dimension, $d^{-1/2}\langle X,Y\rangle=d^{-1/2}\sum_{\ell=1}^{d}X_{\ell}Y_{\ell}$ has mean zero and variance $\sigma_{X}^{2}\sigma_{Y}^{2}$. Furthermore, according to the Lindeberg-Feller Central Limit Theorem, the distribution of the inner product asymptotically converges to the normal distribution as $d$ grows large:
$d^{-1/2}\langle X,Y\rangle\sim\mathcal{N}(0,\sigma_{X}^{2}\sigma_{Y}^{2}).$ (15)
It also follows from the law of large numbers, $d^{-1}\|X\|^{2}=d^{-1}\sum_{\ell=1}^{d}X_{\ell}^{2}$ converges in probability to $\mathbb{E}(X_{\ell}^{2})=\sigma_{X}^{2}$. Similarly $d^{-1}\|Y\|^{2}\to\sigma_{Y}^{2}$ in probability. Therefore,
$\sqrt{d}\cos(X,Y)=\frac{d^{-1/2}\langle X,Y\rangle}{\sqrt{d^{-1}\|X\|^{2}}\sqrt{d^{-1}\|Y\|^{2}}}$
converges to $d^{-1/2}\langle X,Y\rangle/\sigma_{X}\sigma_{Y}$ in probability, and thus (15) gives (14).
Appendix D Details of dimensionality reduction experiments in Section 5
@@
Table 3: The performance of dimensionality reduction for $p$-dimensional embeddings. Rand. stands for Random Order, Skew. for Skewness Sort, and Tour. for Axis Tour. The values in the table correspond to top 1 accuracy for analogy tasks, Spearman’s rank correlation for word similarity tasks, and purity for categorization tasks. Note that at $p=300$, all embeddings give the same results.
Figure 7: The performance of dimensionality reduction for the PCA-transformed embeddings and the Axis Tour embeddings with $\alpha=0,1/3,1/2,1$. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks.
D.1 Detailed explanation of each task
@@
The embedding of the $\text{word}_{i}$ is denoted by $\mathbf{y}_{i}\in\mathbb{R}^{d}$. We used the Google Analogy Test Set Mikolov et al. (2013a), which contains 14 types of analogy tasks, and the Microsoft Research Syntactic Analogies Dataset Mikolov et al. (2013c), which contains 16 types of analogy tasks. In the analogy tasks, the quality of the embeddings is evaluated by inferring $\text{word}_{4}$ to which $\text{word}_{3}$ corresponds if $\text{word}_{1}$ corresponds to $\text{word}_{2}$. We compute the vector $\mathbf{y}_{2}-\mathbf{y}_{1}+\mathbf{y}_{3}$ and see if the closest embedding is $\mathbf{y}_{4}$ (top1 accuracy).
@@ D.3 Setting of $p=d\,(=300)$
Note that the experimental results are the same for all embeddings for $p=300$. First, as we saw in Appendix B.1, when $p=d\,(=300$), the matrix $\mathbf{TF}$ (i.e., the projected $p$-dimensional embeddings) is equal to the matrix $\mathbf{T}$ (i.e., the $d$-dimensional Axis Tour embeddings). Then, by definition, Axis Tour, Random Order, and Skewness Sort are the embeddings obtained by reordering the axes of the ICA-transformed embeddings and flipping their signs as needed. Thus, these three can be seen as the embeddings obtained by applying an orthogonal matrix to the ICA-transformed embeddings. Since the ICA-transformed embeddings are derived from the PCA-transformed embeddings by applying an orthogonal matrix [^17] Refer to the previous work for the relationship between PCA and ICA Yamagiwa et al. (2023)., cosine similarity and Euclidean distance remain unchanged for PCA, Random Order, Skewness Sort, and Axis Tour, leading to identical results in downstream tasks.
Appendix E Additional experiments
E.1 Qualitative observations for all axes of the Axis Tour embeddings
0 1 2 3 4 5 6 7 8 9
phaen region mountain stage italy juan da state india vaas
sandretto goriška mount vinokourov italian spain são terengganu indian jayasuriya
nakhchivan languedoc mountains vuelta di luis paulo kedah singh wicket
burghardt regions everest stages francesco gonzález joão perlis shri jayawardene
regno saguenay-lac-saint-jean peaks magicians pietro garcía janeiro kelantan delhi kallis
10 11 12 13 14 15 16 17 18 19
andrew chief general bhavsar spearritt contributed by ***.com micro-history family
divoff executive then-attorney beppe 297.00 sebti sivuyile tburr 1977-2010 lythraceae
vanwyngarden ceo gen. ayme moçambique avet mulyanda fluto 1951-1972 chrysobalanaceae
hruska justiceship sindiso kahrd epoca aygin boonradom ***.com qc8 polyporaceae
hampsten hienonen jiyane ripstein huajun toosi tolkun ***.com 1977-2006 pyronemataceae
20 21 22 23 24 25 26 27 28 29
order the hungarians serb russian czech germany france canada australia
neuroptera macdougalls ethnic bosnian russia prague german french canadian australian
boletales powhatans asians croatia moscow poland berlin le ontario queensland
svu andhras tatars croatian sergei polish von paris quebec brisbane
poverelle sasanians berbers serbian aleksandr warsaw cologne du saskatchewan perth
30 31 32 33 34 35 36 37 38 39
wiltshire liga coach frank tony mike michael joseph jack &amp;
shrewsbury relegated coaching capra ianno johanns finnissy joe palance llp
lincolnshire fc scolari sinatra canzoneri fetters cerveris macenka o’lantern amp
peterborough f.c. vogts wisner oursler petke stuhlbarg papp spring-heeled &
croydon serie capello aigbogun kornheiser lupica tomasky nollekens lemmon firm
40 41 42 43 44 45 46 47 48 49
analyst an three-week festivities 27th striker goalkeeper mickelson 6-4 2-9
strategist rohp 20-minute celebration 22nd equalized keeper furyk 7-5 8-8
securities average-sized two-week celebrations 26th midfielder goalie els roddick 1-9
hyoty waidhofen two-month commemoration 23rd header goaltender faldo kafelnikov 4-12
udomsirikul taisce 15-minute ceremonies 35th equalised jaaskelainen woods dementieva 17-65
50 51 52 53 54 55 56 57 58 59
24-14 touchdown rebounds inning .301 9-for-13 on island greek george
24-17 quarterback hardaway hitter .292 3-for-8 premised islands greece takei
14-10 touchdowns pippen pitcher .293 5-for-12 picturized fuerteventura athens stroumboulopoulos
20-10 qb shaquille outfielder .289 4-for-11 ixnay conanicut greeks w.bush
27-17 interceptions mcdyess baseman .288 12-of-19 gorging wangerooge zeus maharis
60 61 62 63 64 65 66 67 68 69
president burundi envoys between brouhaha challenges numerous mistake horrific vitriolic
teburoro uganda talks relationship over dilemmas various error terrible racist
vice tanzania envoy quarrel editorship confronting sundry mistakes horrendous denouncing
bagbandy kenya annan rivalry dispute dilemma other misjudgments horrible strident
issayas zambia solana relationships damocles vexing mishaps disregard unspeakable insults
70 71 72 73 74 75 76 77 78 79
disgust tenacity informs life their arguably livelier conclusive endear slowing
anger humility destroys expectancies our liveliest nastier corroboration semblance slowdown
feelings toughness sustains commuted forbears costliest rougher no wobbled weakening
sadness newfound confronts transience fellow-citizens quietest prettier substantiation shaky sluggish
revulsion professionalism goes great-west its best-preserved deadlier scant unscathed decline
80 81 82 83 84 85 86 87 88 89
index 1.48 20.45 30-year 2 sauce beer antidepressant proteins chlorine
asx 1.62 15.55 yield 1 cooked drink drugs protein solvents
ase 2.07 33.65 bond teaspoons cheese drinks drug genes ammonia
ftse 2.05 13.45 10-year 3 roasted brewed medications gene liquid
klse 1.67 17.35 yields teaspoon bread drinkers prozac rna flammable
90 91 92 93 94 95 96 97 98 99
copper oil power line road three-story hotel stores industry subsidiary
nickel petroleum electricity railway highway brick hotels store export company
manganese oilfields substation lines route facade resort grocery manufacturing maker
molybdenum crude hydroelectric arbatsko-pokrovskaya north-south two-story resorts supermarkets distributors alcatel
ore pemex megawatts kolsås two-lane building marriott retailer shoemaking corp.
100 101 102 103 104 105 106 107 108 109
hedge pay land laws court lawsuits charges camp corpses remain
fund fees property regulations judge lawsuit alleged prison corpse remained
funds payments lands enacted appellate litigation prosecutors buchenwald exhumed stayed
investments payment estate law appeals suits indicted camps dismembered stubbornly
investing paid bergisches provisions supreme suit convicted inmates bodies stays
110 111 112 113 114 115 116 117 118 119
hopelessly incredibly ingenious to fostering tricking into has 19,583 3,048
frustrated amazingly devious intend initiatives busily morphed had 21,563 dolne
woefully extremely clever able sustainable concentrating transmuted beeen 16,875 prateek
hamstrung very intricate humiliate empowering classifying degenerates consistantly 20,833 bugis
chronically wonderfully cunning try entrepreneurship disposing delving reinvented 30,313 gumbinnen
Table 4: The top five words for the 00th axis to the $119$th axis when we apply Axis Tour to 300-dimensional GloVe.
120 121 122 123 124 125 126 127 128 129
system desktop phone web television 2300 4:23 _ new 121.58
systems macintosh cellular sites tv 12:30 2:47 ondeck york 114.78
renin-angiotensin pc cellphone site channel 11:30 2:33 holimont orleanians 121.22
centralized software telephone online broadcast 0330 8:38 hawksnest orleans-based 119.76
computerised server wireless myspace cbs 0930 2:21 belleayre n.y. 106.12
130 131 132 133 134 135 136 137 138 139
seventies 2001-2003 four 89 445 1,149 209.6 95.3 28.4 5.5
sixties 2003-2005 five 83 244 4,461 218.8 96.2 23.1 4.8
2010s 1999-2001 three 86 285 1,737 218.3 89.8 23.9 5.7
1800s 1998-2001 eleven 85 292 1,701 308.9 89.9 26.1 4.4
early 1995-1997 six 88 344 1,109 246.3 93.1 27.6 4.6
140 141 142 143 144 145 146 147 148 149
23,000 union alike theories church sex actress german-born old multibillion
110,000 workers academics theory anglican sexual wife newspaperman 35-year 20-million
39,000 machinists industrialists posited episcopal homosexual mother entrepreneur 14-year multi-million
43,000 unions sociologists notions congregations heterosexual daughter politician 50-year 100-million
48,000 teamsters pundits notion congregation unmarried née russian-born 22-year multimillion
150 151 152 153 154 155 156 157 158 159
) rights group force insurgents guns wheels sedan schumacher preakness
( human groups gendarmerie militants rifles wheel sedans barrichello belmont
.0358 aprodeh lashkar-e-jhangvi contingent gunmen gun rudder v8 massa filly
3.7996 zimrights forzani 500-strong guerrillas caliber brakes turbocharged ferrari baffert
unitals pillay harkat contingents forces weapon hydraulic camry raikkonen thoroughbred
160 161 162 163 164 165 166 167 168 169
400m sports welterweight la que basidiomycota .2667 lb3 drawno behshahr
100m sport heavyweight cerva pero gorlice estrategia 5lb krośniewice abyek
200m softball middleweight cenerentola sus empleos creado 8lb brochów ramian
freestyle volleyball ibf ferrière en autovía hirta lb7 lubaczów khamir
200-meter soccer wba louvière una pudiera a.m.-6 4lb pobiedziska javanrud
170 171 172 173 174 175 176 177 178 179
bank river lake city international thailand minister spokesman nafez adtac
central ljubljanica mývatn ozamis airport thai prime faizasyah sedwill coronae
nivard tigris erhai malaybalay tocumen nakhon minster cirtek british-controlled maḩalleh
banks tributary chilka phenix gökçen bangkok minsiter tsanchev videoton decarboxylase
pboc rivers waramaug hitec aiport thais interior ladsous 2-6-2 poderoso
180 181 182 183 184 185 186 187 188 189
8a-4p qeshlaqi automoviles jreyes ***-***-**** _____… moodin - :
romik dizaj-e vx-6 ilovar .0210 interbk ***-***-**** skway bkh www.***.com
shefer kalayeh ***-***-**** hohtz ***-***-**** ooooo… wehz thoh wc2006-asia http://www.***.org
samayoa now-e principalmente ray-finned ***@***.com ***-***-**** prahnk kursh manutd http://www.***.com
hermanas patryan yengejeh odalovic 65stk harrynytimes eesah tsih pickup5thgraf http://www.***.org
190 191 192 193 194 195 196 197 198 199
; sarā rd2 ipnotinmx gibbosa 1507.50 ixmiquilpan prolegs masuku artayev
priu as clientes mordella pratylenchus .000663 ***@***.com grij quartic autuori
chahār weinzapfel significado bifrenaria analista .71078 macul havner tayyab
.8226 khvosh desempleo sandwicense fimbriatus ramnarine jagdeep rambusch shealy pareto
sa/b akbarābād gevar ***-***-**** laticeps hàm naab yr.ago vidro zermelo
200 201 202 203 204 205 206 207 208 209
swaffham phthalic 72-77 magnifica funderburke cunxiao wang japan south park
**-***-**** chamba gildernew compra hintikka cunshen zhang japanese korean prenton
ollivier nguyên deanne qc7 chofetz cunxu xu tokyo korea ji-sung
zend kinghorn ethelbert kowalska danys siyuan li akira dakotans naturpark
cronulla-sutherland matanuska-susitna kether handanovic opticians lanqing liu takashi carolinians breffni
210 211 212 213 214 215 216 217 218 219
national parliament party polling republican . i silently trousers whitish
winema duma janata electoral sen. , ’cause stared pants yellowish
jurnalul speaker bjp balloting gop but gonna screamed dresses brownish
ranthambhore 150-seat socialists election republicans though yeah yelled dress greyish
chúa 500-seat ndp tallying mccain even gosh sobbing wearing reddish-brown
220 221 222 223 224 225 226 227 228 229
ultraviolet spacecraft aircraft vessel hurricane temperatures grasslands shrubs mammals virus
infrared astronauts jet ship storm humid habitats trees birds h5n1
telescopes astronaut planes ships storms chilly soils vines rabbits swine
light nasa boeing vessels typhoon weather marshes planted animals flu
wavelengths orbit f-15 boats cyclone unseasonably sediments seedlings reptiles outbreak
230 231 232 233 234 235 236 237 238 239
diagnosed medical undergo survey agency newspaper ign award film superhero
lung care thorough surveys notimex daily popmatters awards films marvel
inflammation hospital undergoing statistics kathpress zeitung allmusic awarded movie spin-off
complications hospitals evaluation gallup bss izvestia reviewer prize starring superheroes
fractures physician undergone statistical telam kommersant gamespot emmy directed characters
Table 5: The top five words for the $120$th axis to the $239$th axis when we apply Axis Tour to 300-dimensional GloVe.
240 241 242 243 244 245 246 247 248 249
album piano paintings manuscript language name formula_1 set mark brian
albums violin painting biographies languages names formula_2 aflame high-water trenchard-smith
band cello art pages pashto surname formula_3 10-cd dindal doyle-murray
self-titled percussion sculpture book colloquial phrase formula_4 setting wainberg cadd
ep orchestral watercolor handwritten dialect misspelling formula_5 4-cd hoppus donlevy
250 251 252 253 254 255 256 257 258 259
ireland martin peter paul john bazzani david scott 2-54
belfast scelzo guillam virilio rhys-davies looking-glass pittu livengood cretier
irish wansleben shockheaded ricoeur canemaker xff mouret proval lobdell veltman
mowlam clunes maffay langmack motson munro margolick gud f.r.
northern damm rauhofer celan mcgahern schizopolis mármol gorcyca speedman nahant
260 261 262 263 264 265 266 267 268 269
palm county calif. school university 1976 1853 february king alwaleed
beach unincorporated california high professor 1973 1852 october sigismund prince
fla. dekalb inglewood elementary graduate 1966 1847 december iv saudi
ostrowski fayette pasadena pine-richland doctorate 1968 1856 june emperor kingdom
broward pulaski pomona jr/sr faculty 1971 1854 april iii tupouto’a
270 271 272 273 274 275 276 277 278 279
ali israel daniel swedish thomas william charles james wilkison nack
al israeli briere sweden kretschmann mastrosimone dutoit frain mudavadi narrowly
mohammed israelis macivor norwegian aquinas beaudine grodin roday henwood gyoergy
sheikh netanyahu gildenlöw norway quasthoff fichtner sheeler luceno hogzilla khand
mohammad aviv balavoine fredrik fingar saroyan wuorinen remar mottola vanak
280 281 282 283 284 285 286 287 288 289
nhls janyk vranjes robert van chris marchena steve shaara richard
ba’asyir tookie kilvert deniro lieshout volstad staniforth railsback jeff stoltzman
organizaciones germanica bukan halmi dijk braide kaboul turre shain ayoade
yordanov walthers paratore garrigus zandt bordano pelzer forbert friesen basehart
millán adquisiciones ponge redford tuyl tidland 43.18 kimock sharlet sandomir
290 291 292 293 294 295 296 297 298 299
simon malatesta bilyaletdinov .000106 nazione olya gearon sirajul wonk kerberos
boccanegra .0217 pouget fhs ba872 dodecahedral mikuláš overexcited kappa mckelvey
napier-bell lockard cytidine colecovision blouin cih bracigliano nabhani fraternity wajir
vouet 330-pound zhulali ronghua comunicacion alnus venero then-reigning godsmack veg
callow puji bc4 yunlong wachau 9.29 mariangela karmichael bibliophiles pdca
Table 6: The top five words for the $240$th axis to the $299$th axis when we apply Axis Tour to 300-dimensional GloVe.
In Table 1, we used 300-dimensional GloVe and showed the semantic continuity of the axes of the Axis Tour embeddings, with illustrative examples of the meaning of countries (the $23$rd axis to the $31$st axis), law (the $101$st axis to the $109$th axis), and art (the $237$th axis to the $245$th axis). This section presents the top five words of the normalized embeddings across all 300 axes in Tables 45,and 6.
For example, in Table 4, the $45$th and $46$th axes are related to soccer, the $47$th axis to golf, the $48$th axis to tennis, the $49$th and $50$th axes to scores, the $51$st axis to American football, the $52$nd axis to basketball, and the $53$rd axis to baseball. These axes illustrate the semantic continuity across sports.
In Table 5, the $129$th to the $140$th axes are related to numbers. It is also interesting to note that the top words of each axis have similar numerical scales, highlighting how well the axis of the ICA-transformed embeddings captures meaning. In addition, the meaning of each axis changes continuously, much like a game of word association: the $219$th axis relates to colors, the 220th axis to light, the 221st axis to space, the 222nd axis to airplanes, the 223rd axis to ships, the 224th axis to storms, the 225th axis to weather, the 226th axis to biomes, the 227th axis to plants.
In Table 6, the 272nd to the 289th axes are related to personal names from different linguistic regions as if this were a cluster of the meaning.
Note that due to space limitations, the top 1 and top 3 words on the $185$th axis are truncated because they are repetitive symbols, and that URLs, email addresses, and phone numbers are anonymized.
E.2 Comparison of $\alpha$